# MetaWFN: A Platform for Unified Implementation of Many-Electron Wavefunctions

### v1

@@ 3.2.3 Integral Manager
The IntegralManager class is responsible for handling molecular integrals. The main challenge is how to deal with different kinds of permutation and single/double point group symmetries. In the spin-free case, the two-electron integrals exhibit 8-fold permutation symmetry and are real; for systems with $L_{z}$ symmetry, this symmetry is reduced to 4-fold, but the integrals still remain real. In the spin-dependent case, the integrals (over spinors) are generally complex with complicated permutation and double point group symmetries7. Hence, permutation and single/double point group symmetries are not uncorrelated. To store and retrieve different kinds of molecular integrals in a uniform manner, we used a tag system to label different permutation symmetries:
@@
Here, we use a metafunction perm_symm_t to map each single/double group irrep to the corresponding permutation symmetry for molecular integrals at compile time. The skeleton code for the IntegralManager class is:
@@
8virtual ScalarTy get_1b(const unsigned p, const unsigned q) const
9{
@@
11}
12virtual ScalarTy get_2b(const unsigned p, const unsigned q, const unsigned r, const unsigned s) const
13{
@@
23VectorStorage<ScalarTy, perm_symm_t<SymTy>> integrals_1b;
24VectorStorage<ScalarTy, perm_symm_t<SymTy>> integrals_2b;
@@
The template parameters ScalarTy and perm_symm_t<SymTy> of VectorStorage specify scalar type (real or complex) and permutation symmetry of the integrals, respectively. Note in passing that point group symmetries are not employed here for storage reduction, simply because all integrals can be stored in memory after accounting for permutation symmetries (for the number of correlated orbitals is usually much less 1000). Yet, point group symmetries are indeed ensured by the get_1b and get_2b functions, which return zeros when the direct product of the irreps of the input orbitals does not contain the trivial irrep. With this design, the C++ compiler can use the tagging and type-trait system to synthesize the correct data structures for storing the molecular integrals at compile time. This allows the programmer to write a uniform code without the need to consider the permutation symmetry explicitly.
@@ 3.2.4 Evaluation of HMEs
In MetaWFN, the calculation of HMEs involves several steps. The general workflow begins with the classification of CFG pairs and generation of the corresponding reduced CFG pairs. This is followed by retrieving the necessary integrals, computing the BCCs, and finally contracting these components to obtain the HMEs. Due to the complexity and intricacy of the algorithms involved in these steps, we do not delve into details here. Instead, we emphasize that MetaWFN provides kernel-level template functions for each step, which can be utilized by higher-level algorithms. For example, for singly excited CFG pairs, we have
@@
This modular design allows for flexibility and reusability, making it easier to maintain and extend the codebase.
An important aspect of our implementation is the accommodation of different Hamiltonians. Different Hamiltonians correspond to distinct molecular integrals, which have different permutation symmetries, and therefore different storage requirements, necessitating specialized implementations for each. To handle this variety without code duplication, MetaWFN employs a technique known as tag dispatching, where specific types (tags) are used to direct the compiler to appropriate function implementations at compile time via function overloading. For instance, the last parameter in the functions _generate_table_sngl and _fetch_ints_sngl is a tag that specifies the type of Hamiltonian, which allows the compiler to pick up automatically the right Hamiltonian.
3.3 Algorithm Layer
The algorithm layer is the highest level in the MetaWFN architecture. It employes the classes defined in the quantum chemistry layer as template parameters to generate the corresponding algorithm modules, such as selection and perturbation (see Sec. sCIPT2). These algorithm classes follow a consistent pattern: First, each algorithm is encapsulated in a class (typically ending with ThreadContext) that describes the tasks to be performed by a single thread. For instance, in the ranking step of the selection process, the RankingThreadContext class is invoked to generate important excited CSFs/DETs from a reference CSF/DET. In the PT2 module, a similar class handles the calculation of the PT2 correction from a specific subspace $W_{i}$. Second, parallelization is achieved through parallel templates that describe how to distribute and manage tasks across multiple threads or processes. For OpenMP, these algorithm classes inherit from a base class (e.g., OpenMPAlgorithmBase described above) and implement the necessary functions to assign tasks to each thread and merge the results efficiently. An important aspect of our design is the implementation of MPI: The same OpenMP parallel algorithm classes are inherited, with suitable modifications of the relevant functions to create an automatic mapping functor from OpenMP to MPI. Since the sCI algorithms can be designed such that subtasks are independent of each other and require no communication during execution—only data merging in the end, this functor can effectively distribute tasks across MPI processes. Separating MPI from the single-threaded algorithms ensures that the core algorithm logic remains unchanged while parallelism is handled independently. Specific implementation of the algorithm layer is skipped here, as it would largely repeat the content already covered in Sec. sCIPT2.
3.4 Overview of MetaWFN Functionality
MetaWFN offers a comprehensive suite of features designed to facilitate the development of quantum chemical methods. The key functionalities currently supported by MetaWFN include:
@@ 3.4 Overview of MetaWFN Functionality
    • Non-relativistic Hamiltonian: iCIPT213, 14
@@
    • spin-separated X2C Hamiltonian64, 65: SOiCI and iCISO33
@@ 4 Application
MetaWFN has served as the platform for the development of spin-free13, 14 and spin-dependent33, 7 iCIPT2, which can be applied to both core and valence states of challenging systems, especially for benchmarking purposes69, 70, 71. Taking iCIPT2 as the solver of CASSCF, we obtain iCISCF(2)15 and SOiCISCF(2)34, where the symbol ‘(2)’ implies that the PT2 corrections are only performed within the active space. iCISCF(2) can handle active spaces as large as (60e,60o), on top of which spin-orbit coupling can further be treated variationally by SOiCISCF(2).
@@
In principle, the costs of the ranking, pruning, and PT2 steps of the three methods should be compared separately. However, the costs for the ranking and pruning steps exhibit considerable variations due to several factors: (1) the number of iterations, including the macro-cycles of ranking and pruning as well as the micro-cycles of diagonalization, may differ significantly; (2) the ranking step is governed predominantly by non-computational operations such as memory allocation and sorting; (3) even the computational operations are memory-bound (meaning that their performance depends on data retrieval from cache/memory) and cannot fully explore the data locality in practice. Consequently, the running environment of the computer system may impact significantly the performance of these operations. As a result, it not meaningful to compare the costs for the ranking and pruning steps of different methods. In contrast, the PT2 step is much less affected by environmental factors. In particular, it dominates the overall computational cost as the size of the variational space grows. Therefore, it is more appropriate to compare the PT2 wall times ($T_{2}$) in the sf-X2C-iCIPT, SOiCI, and 4C-iCIPT2 calculations. The results are documented in Table 3, along with the sizes $|P|$ of the variational spaces for the selection thresholds $C_{\text{min}}=\{5.0,3.0,1.5,0.9,0.5\}\times 10^{-5}$. It is first noted that the PT2 steps of all the methods scale linearly with respect to the sizes of the variational spaces (cf. Table 4). The PT2 step of SOiCI turns out be about $1.5\sim 2$ times that of sf-X2C-iCIPT2. This can be understood as follows. If the additional selection of singly excited CSFs by the effective one-body spin-orbit coupling (SOC) operator (so-DKH164, 65) is not to be performed after the selection of sf-X2C-iCIPT2, the PT2 step of SOiCI would precisely double that of sf-X2C-iCIPT2, due to the doubled size of the variational space (i.e., $22S+1=$ 2 times $|P|$ of sf-X2C-iCIPT2). In the case of Br, only a small number of quartet CSFs is brought into the variational space of SOiCI due to the additional SOC selection. Therefore, the fact that the PT2 step of SOiCI can be less than double that of sf-X2C-iCIPT2 must stem from the better data vectorization in the case of SOiCI, which treats the three states simultaneously rather than individually as done by sf-X2C-iCIPT2. In contrast, the PT2 step of 4C-iCIPT2 is typically $10\sim 15$ times that of SOiCI, due to changes in the natures of MPBFs and Hamiltonian.
@@
    Spin-orbit splitting in cm-1 (experimental value74: 3685 cm-1).
@@ 5 Conclusion and Outlook
By utilizing advanced C++ features, such as template metaprogramming and polymorphism, a highly modular, extendable, and efficient platform, MetaWFN, has been introduced for unified implementation of many-electron wavefunctions. It is composed of three abstract layers, i.e., infrastructure, quantum chemistry, and algorithm. The infrastructure layer supplies unified interfaces for basic linear algebra operations and algorithm templates for parallel computing, abstracting scalar type dependencies and parallel strategies. The quantum chemistry layer addresses domain-specific components such as Hamiltonians, MPBFs, symmetries, integrals, and matrix elements, employing a type-trait and tagging system. The algorithm layer implements high-level computational algorithms, benefiting from the abstractions provided by the infrastructure and quantum chemistry layers. This way, wavefunction methods, quantum chemistry information, and parallel strategies are decoupled to the maximal extent, thereby reducing code repetition and enhancing code extendability for new functionalities. Among others, some immediate developments include (1) GPU acceleration of both matrix-vector products and perturbation corrections, and (2) support of more and general Hamiltonians so as to extend the applicability of MetaWFN to a wider range of fermionic and/or bosonic quantum systems.

### v2

@@ 3.2.3 Integral Manager
The IntegralManager class is responsible for handling molecular integrals. The main challenge is how to deal with different kinds of permutation and single/double point group symmetries. In the spin-free case, the two-electron integrals exhibit 8-fold permutation symmetry and are real; for systems with $L_{z}$ symmetry, this symmetry is reduced to 4-fold, but the integrals still remain real. In the spin-dependent case, the integrals (over spinors) are generally complex with complicated permutation and double point group symmetries7. Hence, permutation and single/double point group symmetries are not uncorrelated. To store and retrieve different kinds of molecular integrals in a uniform manner, we used a tag system to label different permutation symmetries (Listing 12) and a metafunction perm_symm_t to map each single/double group irrep to the corresponding permutation symmetry for molecular integrals at compile time.
Listing 12: Tag system to label permutation symmetries.
@@
The skeleton code for the IntegralManager class is shown in Listing 13. The template parameters ScalarTy and perm_symm_t<SymTy> of VectorStorage specify scalar type (real or complex) and permutation symmetry of the integrals, respectively. Note in passing that point group symmetries are not employed here for storage reduction, simply because all integrals can be stored in memory after accounting for permutation symmetries (for the number of correlated orbitals is usually much less 1000). Yet, point group symmetries are indeed ensured by the get_1b and get_2b functions, which return zeros when the direct product of the irreps of the input orbitals does not contain the trivial irrep. With this design, the C++ compiler can use the tagging and type-trait system to synthesize the correct data structures for storing the molecular integrals at compile time. This allows the programmer to write a uniform code without the need to consider the permutation symmetry explicitly.
Listing 13: Skeleton of class IntegralManager.
@@
8 virtual ScalarTy get_1b(const unsigned p, const unsigned q) const
9 {
@@
11 }
12 virtual ScalarTy get_2b(const unsigned p, const unsigned q, const unsigned r, const unsigned s) const
13 {
@@
23 VectorStorage<ScalarTy, perm_symm_t<SymTy>> integrals_1b;
24 VectorStorage<ScalarTy, perm_symm_t<SymTy>> integrals_2b;
@@ 3.2.4 Evaluation of HMEs
In MetaWave, the calculation of HMEs involves several steps. The general workflow begins with the classification of oCFG pairs and generation of the corresponding reduced oCFG pairs. This is followed by retrieving the necessary integrals, computing the BCCs, and finally contracting these components to obtain the HMEs. Due to the complexity and intricacy of the algorithms involved in these steps, we do not delve into details here. Instead, we emphasize that MetaWave provides kernel-level template functions for each step, which can be utilized by higher-level algorithms. For example, for singly excited oCFG pairs, we have several functions shown in Listing 14. This modular design allows for flexibility and reusability, making it easier to maintain and extend the codebase.
Listing 14: Functions for evaluation of HMEs between singly excited oCFG pairs.
@@
An important aspect of our implementation is the accommodation of different Hamiltonians. Different Hamiltonians correspond to distinct molecular integrals, which have different permutation symmetries, and therefore different storage requirements, necessitating specialized implementations for each. To handle this variety without code duplication, MetaWave employs a technique known as tag dispatching, where specific types (tags) are used to direct the compiler to appropriate function implementations at compile time via function overloading. For instance, the last parameter in the functions _generate_table_sngl and _fetch_ints_sngl is a tag that specifies the type of Hamiltonian, which allows the compiler to pick up automatically the right Hamiltonian.
3.3 Wavefunction Layer
The wavefunction layer is the highest level in the MetaWave architecture. It employes the classes defined in the Hamiltonian layer as template parameters to generate the corresponding algorithm modules, such as selection and perturbation (see Sec. 2). These algorithm classes follow a consistent pattern: First, each algorithm is encapsulated in a class (typically ending with ThreadContext) that describes the tasks to be performed by a single thread. For instance, in the ranking step of the selection process, the RankingThreadContext class is invoked to generate important excited CSFs/DETs from a reference CSF/DET. In the PT2 module, a similar class handles the calculation of the PT2 correction from a specific subspace $W_{i}$. Second, parallelization is achieved through parallel templates that describe how to distribute and manage tasks across multiple threads or processes. For OpenMP, these algorithm classes inherit from a base class (e.g., OpenMPAlgorithmBase described above) and implement the necessary functions to assign tasks to each thread and merge the results efficiently. An important aspect of our design is the implementation of MPI: The same OpenMP parallel algorithm classes are inherited, with suitable modifications of the relevant functions to create an automatic mapping functor from OpenMP to MPI. Since the sCI algorithms can be designed such that subtasks are independent of each other and require no communication (aside from dynamic scheduling) during execution—only data merging in the end, this functor can effectively distribute tasks across MPI processes. Separating MPI from the single-threaded algorithms ensures that the core algorithm logic remains unchanged while parallelism is handled independently. Specific implementation of the wavefunction layer is skipped here, as it would largely repeat the content already covered in Sec. 2.
3.4 Overview of MetaWave Functionality
MetaWave offers a comprehensive suite of features designed to facilitate the development of quantum chemical methods. The key functionalities currently supported by MetaWave include:
@@ 3.4 Overview of MetaWave Functionality
    • Nonrelativistic Hamiltonian: iCIPT216, 17
@@
    • spin-separated X2C Hamiltonian73, 74: SOiCI and iCISO36
@@ 4 Application
MetaWave has served as the platform for the development of spin-free16, 17 and spin-dependent36, 7 iCIPT2, which can be applied to both core and valence states of challenging systems, especially for benchmarking purposes81, 82, 83. Taking iCIPT2 as the solver of CASSCF, we obtain iCISCF(2)18 and SOiCISCF(2)37, where the symbol ‘(2)’ implies that the PT2 corrections are only performed within the active space. iCISCF(2) can handle active spaces as large as (60e,60o), on top of which spin-orbit coupling can further be treated variationally by SOiCISCF(2).
@@
In principle, the costs of the ranking, pruning, and PT2 steps of the three methods should be compared separately. However, the costs for the ranking and pruning steps exhibit considerable variations due to several factors: (1) the number of iterations, including the macro-cycles of ranking and pruning as well as the micro-cycles of diagonalization, may differ significantly; (2) the ranking step is governed predominantly by non-computational operations such as memory allocation and sorting; (3) even the computational operations are memory-bound (meaning that their performance depends on data retrieval from cache/memory) and cannot fully explore the data locality in practice. Consequently, the running environment of the computer system may impact significantly the performance of these operations. As a result, it is not meaningful to compare the costs for the ranking and pruning steps of different methods. In contrast, the PT2 step is much less affected by environmental factors. In particular, it dominates the overall computational cost as the size of the variational space grows. Therefore, it is more appropriate to compare the PT2 wall times ($T_{2}$) in the sf-X2C-iCIPT, SOiCI, and 4C-iCIPT2 calculations. The results are documented in Table 3, along with the sizes $|P|$ of the variational spaces for the selection thresholds $C_{\text{min}}=\{5.0,3.0,1.5,0.9,0.5\}\times 10^{-5}$. It is first noted that the PT2 steps of all the methods scale linearly with respect to the sizes of the variational spaces (cf. Table 4). The PT2 step of SOiCI turns out be about $1.5\sim 2$ times that of sf-X2C-iCIPT2. This can be understood as follows. If the additional selection of singly excited CSFs by the effective one-body spin-orbit coupling (SOC) operator (so-DKH173, 74) is not to be performed after the selection of sf-X2C-iCIPT2, the PT2 step of SOiCI would precisely double that of sf-X2C-iCIPT2, due to the doubled size of the variational space (i.e., $22S+1=$ 2 times $|P|$ of sf-X2C-iCIPT2). In the case of Br, only a small number of quartet CSFs is brought into the variational space of SOiCI due to the additional SOC selection. Therefore, the fact that the PT2 step of SOiCI can be less than double that of sf-X2C-iCIPT2 must stem from the better data vectorization in the case of SOiCI, which treats the three states simultaneously rather than individually as done by sf-X2C-iCIPT2. In contrast, the PT2 step of 4C-iCIPT2 is typically $10\sim 15$ times that of SOiCI, due to changes in the natures of MPBFs and Hamiltonian.
@@
    Spin-orbit splitting in cm-1 (experimental value86: 3685 cm-1).
@@ 5 Conclusion and Outlook
By utilizing advanced C++ features, such as template metaprogramming and polymorphism, a highly modular, extendable, and efficient platform, dubbed MetaWave, has been introduced for unified implementation of many-electron wavefunctions. It is composed of three abstract layers, i.e., infrastructure, Hamiltonian, and wavefunction. The infrastructure layer supplies unified interfaces for basic linear algebra operations and algorithm templates for parallel computing, abstracting scalar type dependencies and parallel strategies. The Hamiltonian layer addresses domain-specific components such as Hamiltonians, MPBFs, symmetries, integrals, and matrix elements, employing a type-trait and tagging system. The wavefunction layer implements high-level computational algorithms, benefiting from the abstractions provided by the infrastructure and Hamiltonian layers. In this way, Hamiltonians, wavefunctions, and parallel strategies are decoupled to the maximal extent, thereby reducing code repetition and enhancing code extendability for new functionalities. Among others, some immediate developments include (1) GPU acceleration of both matrix-vector products and perturbation corrections, (2) support of more and general Hamiltonians for the extension of MetaWave to a wider range of fermionic and/or bosonic quantum systems, and (3) support of tensor contractions for unified implementation of relativistic and nonrelativistic coupled-cluster methods.